# HKCanto-Eval: A Benchmark for Evaluating Cantonese Language Understanding and Cultural Comprehension in LLMs

### v1

@@ 4.4 Linguistic and NLP Tasks
Most LLMs understand Cantonese fine. Most models performed well in Sentiment Analysis (GPT4o 79.7%, Llama 3.1 405B 78.8%), Translation (3-shot: GPT4o 98.3%, Qwen2 72B 96.6%), and Summarisation (Claude 3.5 Sonnet 92.7%, Gemma 2 9B 91.3%). Models that obtained lower scores are often due to task completion problems, e.g. failure to handle long input and problems with low frequency/mixed-language tokens.
Proprietary and large open-weight models have good Cantonese lexical knowledge. The performance in translation and sentiment analysis is closely tied to the ability to determine the meaning of Cantonese-specific words that are not found or used differently in Mandarin. Most models also performed well in the Character Selection sub-task under Orthographic Knowledge. It is noteworthy that despite good performance with proprietary models (73.1% - 88.5%) and some open-weight models (CLLM 34B and Qwen2 72B both 76.9%), GPT4 struggled with Cantonese orthography (65.4%).
LLMs in general lack knowledge about Cantonese pronunciation. Most models failed to conduct simple tasks such as judging homophone or rhyme pairs, with GPT-4o being a notable exception (Homophone: 56.0%; Rhyming: 96.0%). Poor performance in other models is influenced by Mandarin homophones. In the Grapheme-to-Phoneme (G2P) conversion task, all models performed worse than the rule-based control (Visual Fonts, CER 5.0%), the closest being GPT4o (5.4%), Claude 3.5 Sonnet (7.9%). All models performed at chance level in phonological reasoning, which is particularly challenging for Cantonese due to limited specialised data.
LLMs in general do not have meta-linguistic knowledge represented in Cantonese. Although certain models especially the Chinese proprietary models performed well in the visual similarity task (SenseChat 70%, Doubao 70%, Ernie 75%) or orthographic reasoning (GPT4o 63.0%), the knowledge seems to have come from Simplified Chinese, thus their good performance is not transferred to Cantonese-specific items. This seems to be caused by insufficient descriptive knowledge about the structure and properties associated with the individual glyphs.
Model Linguistic NLP
Claude 3.5 Sonnet 70.8% 89.2%
Doubao Pro 64.7% 87.0%
Ernie 4.0 62.8% 82.7%
Gemini 1.5 Flash 55.2% 83.2%
Gemini 1.5 Pro 66.4% 87.9%
GPT4o 76.7% 89.6%
GPT4o-mini 64.7% 86.1%
SenseChat 52.8% 78.8%
Aya 23 8B 19.0% 70.1%
CLLM 6B 17.0% 71.9%
CLLM 34B 59.7% 73.3%
Yi 1.5 6B 21.3% 56.6%
Yi 1.5 9B 28.0% 72.2%
Yi 1.5 34B 25.4% 82.9%
Gemma 2 2B 24.5% 73.4%
Gemma 2 9B 40.3% 85.0%
Gemma 2 27B 42.1% 83.2%
Llama 3.1 8B 39.0% 60.3%
Llama 3.1 70B 56.0% 84.5%
Llama 3.1 405B 60.6% 64.4%
Mistral Nemo 12B 41.5% 68.8%
Qwen2 7B 27.8% 66.8%
Qwen2 72B 43.3% 83.5%
Control 60.3% 76.8%
Table 2: Model performance in linguistics and NLP tasks
@@ 5 Conclusion
This paper presents HKCanto-Eval, the first comprehensive evaluation benchmark focussing on Hong Kong Cantonese, by comparing the Cantonese language support of 6 proprietary and 7 open-weight model families. Our findings indicate that while these models can understand Cantonese in various contexts, retrieve knowledge about Hong Kong, and address problems written in or about Cantonese to some extent, there are notable limitations. Most models, especially open-weight models in the 6–9B range, lack sufficient linguistic, cultural and professional knowledge in Cantonese and Hong Kong. Performance was particularly poor for questions requiring knowledge not commonly found in major online sources.
One area that we paid close attention to is the presence of metalinguistic knowledge in these models. There is concern that models showed Cantonese proficiency in linguistic and NLP tasks primarily through Mandarin. If their linguistic understanding is based solely on Mandarin, they may perform well on simpler tasks but struggle significantly with “false friends" between languages, as Mandarin knowledge becomes a hindrance. This benchmark introduces a novel perspective, focussing on Cantonese processing abilities beyond superficial slang and expressions. By requiring reasoning about sounds and characters specific to Cantonese, our benchmark provides a fairer judgement that credits models accurately capturing Cantonese phonology and orthography, while exposing those that appear competent in Cantonese but are heavily reliant on Mandarin.
@@ 6 Limitations & Future Directions
The current benchmark exhibits several limitations. First, the use of machine translation introduces potential inaccuracies. While Gemini 1.5 Flash balances cost and quality, human-translated questions could provide a more reliable benchmark, albeit at a higher resource cost. The reliance on multiple-choice and text-based questions does not fully capture the capabilities required for practical LLM applications such as code generation and mathematical problem-solving, which demand coherent and contextual text generation. The dataset also lacks multi-modal data like image and audio, which is now supported by proprietary models and should be evaluated.
The newly and manually created questions might contain biases and a lack of scalability and comprehensiveness. The cultural questions, predominantly created by colleagues and relatives of the authors, may introduce bias in cultural references and wordings, leading to an over-representation of certain perspectives while under-representing others, such as traditional practices. Political topics were also specifically excluded, limiting cultural representation.
Human evaluation, while insightful, is not scalable. Automated and objective evaluation methods, such as LLM-as-a-judge or rule-based approaches, are necessary for efficient evaluation, but this is challenging due to the low-resource nature of Cantonese.
@@ Appendix A Questionable Practices in an Existing Work
Jiang et al. (2024) recently proposed a Cantonese evaluation dataset consisting of 5 datasets. The authors cited a HuggingFace organisation homepage in a footnote for their translation dataset but offered no further specifics. Yet, the dataset’s entries were not identifiable within the linked account. Nonetheless, the translation pairs between Cantonese and English are dubious. During the machine translation process, the less advanced models before GPT3.5 tend to break English sentences into phases by punctuation marks or connecting words such as "and" and "but". Then the phases were translated individually and finally put together into one sentence in Cantonese. As a result, the translated texts were full of wrong wordings such as the following translation pair:
@@
The translation service broke down the original English sentences by punctuation and connectives (such as "and"). Then, it translated the broken-down phases individually and joined them into a Cantonese/Chinese sentence, resulting in awkward and incorrect punctuation usage. The heavy use of poor machine translation of existing work offered little in terms of novel contributions or insights into the language.
@@ Appendix C Academic and Professional Dataset
    Insurance Intermediaries Qualifying Examination: The Insurance Authority in Hong Kong regulates and licenses all insurance intermediaries in Hong Kong and offers this exam for any person who wishes to carry out insurance activity.
@@
    Licensing Exam for Securities & Futures Intermediaries (LE): An exam offered by the Hong Kong Securities and Investment Institute and recognised by the Securities and Futures Commission (SFC). It is a professional qualification exam for people who would like to work in the securities and investment industry in Hong Kong.
@@
The Academic Dataset subset was sourced from scanned copies of the HKDSE exam from the last 2 to 4 years. The details can be found in Table 4. The scanned PDFs were processed to extract text with the Gemini 1.5 Pro API. The extracted text was then filtered by the authors to remove questions that required information from other questions. Those requiring references to visual materials like pictures and maps were also removed but reserved for future evaluation of vision-enabled LLMs.
Although only multiple-choice questions were included in the dataset, more complicated questions that require complex reasoning and human judgment, such as Chinese listening and Liberal Studies exams were studied. Initial testing showed proprietary LLMs can easily achieve perfect scores in the Chinese listening exam when given with the transcript, while the length of the transcript exceeded some open source LLMs’ context length. Further experiments were conducted with Google Gemini 1.5 Flash and Pro to take the audio file and written questions as input. Although the two models again achieved a perfect score, no other model supported audio input at the time of the experiment, so the task was excluded. For the now-defunct Liberal Studies, examinees were required to write a passage taking into account the provided reference materials and then develop their own point of view supported by examples from the candidates’ own knowledge. However, initial testing showed that none of the LLMs were able to include their own examples despite explicitly requested in the prompt.
@@ Appendix E Linguistic Knowledge Dataset
Phonological Knowledge Orthographic Knowledge
Model
Homo-
phone
Rhyme Misc. Avg.
Visual
Similarity
Canton.
Char.
Misc. Avg.
Claude 3.5 Sonnet 28.0% 64.0% 16.0% 31.0% 50.0% 76.9% 59.3% 64.5%
Doubao Pro 16.0% 44.0% 16.0% 23.0% 70.0% 80.8% 48.1% 64.5%
Ernie 4.0 28.0% 60.0% 18.0% 31.0% 70.0% 80.8% 53.7% 67.5%
Gemini 1.5 Flash 12.0% 20.0% 24.0% 20.0% 40.0% 73.1% 31.5% 46.0%
Gemini 1.5 Pro 16.0% 40.0% 24.0% 26.0% 50.0% 88.5% 46.3% 60.5%
GPT4o 56.0% 96.0% 28.0% 52.0% 50.0% 65.4% 63.0% 63.5%
GPT4o-mini 20.0% 60.0% 20.0% 30.0% 30.0% 57.7% 40.7% 44.5%
SenseChat 16.0% 36.0% 22.0% 24.0% 75.0% 76.9% 42.6% 61.8%
Aya 23 8B 12.0% 40.0% 14.0% 20.0% 15.0% 19.2% 31.5% 25.8%
CLLM 6B 24.0% 8.0% 18.0% 17.0% 20.0% 50.0% 27.8% 33.0%
CLLM 34B 28.0% 28.0% 14.0% 21.0% 35.0% 76.9% 37.0% 48.8%
Yi 1.5 6B 28.0% 12.0% 12.0% 16.0% 10.0% 50.0% 20.4% 26.5%
Yi 1.5 9B 36.0% 40.0% 24.0% 31.0% 30.0% 57.7% 18.5% 32.5%
Yi 1.5 34B 16.0% 32.0% 26.0% 25.0% 30.0% 61.5% 33.3% 41.5%
Gemma 2 2B 8.0% 24.0% 18.0% 17.0% 25.0% 53.8% 22.2% 32.3%
Gemma 2 9B 20.0% 28.0% 24.0% 24.0% 25.0% 50.0% 33.3% 37.3%
Gemma 2 27B 20.0% 12.0% 16.0% 16.0% 25.0% 65.4% 24.1% 36.3%
Llama 3.1 8B 12.0% 16.0% 18.0% 16.0% 25.0% 42.3% 38.9% 38.3%
Llama 3.1 70B 28.0% 40.0% 12.0% 23.0% 30.0% 61.5% 35.2% 42.5%
Llama 3.1 405B 20.0% 44.0% 18.0% 25.0% 35.0% 65.4% 50.0% 52.8%
Mistral Nemo 12B 12.0% 28.0% 10.0% 15.0% 25.0% 23.1% 37.0% 32.3%
Qwen2 7B 8.0% 40.0% 12.0% 18.0% 35.0% 46.2% 33.3% 38.8%
Qwen2 72B 12.0% 28.0% 16.0% 18.0% 50.0% 76.9% 48.1% 58.5%
Random 16.0% 28.0% 24.0% 23.0% 30.0% 11.5% 27.8% 25.5%
Table 7: Model performance on Linguistic Knowledge Dataset multiple-choice questions
This dataset comprises three sub-datasets with questions carefully designed with expert input from researchers specialising in different areas of Cantonese linguistics. The structure and questions of each dataset are outlined below. The results of all three can be found in Table 7.
@@ E.1 Phonological Knowledge Dataset
    MultiPron Resolution: A word that contains a character that can be pronounced in multiple ways (like the English word read and be pronounced as red and ri:d), and the task is to decide how this character should be pronounced. For example, if the given word is [Uncaptioned image][Uncaptioned image] and the question asks for a homophone for the 2nd character, given the list A [Uncaptioned image] B [Uncaptioned image] C [Uncaptioned image] D [Uncaptioned image], the answer is none of the above. The given character should be pronounced as ho4 and none of the answer (ho2, ho6, ho1, wo4) matches this pronunciation.
@@
    Poetry Rhyme: Given a poem, choose the correct combination of rhyming characters, e.g. in the poem [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]。 The last syllables are zeon6, lau4, muk6, lau4 in Cantonese, [Uncaptioned image][Uncaptioned image] are the two characters that rhyme in the poem when recited in Cantonese.
@@ E.3 Grapheme-to-Phoneme (G2P) Conversion Dataset
Score calculation: All acceptable variants have been listed as answers and all variants are considered equally good. This is to handle variant forms and ambiguous interpretations that may not be the standard, but native speakers of Cantonese accept. We will use the answer with the highest score for the subsequent calculation. Two metrics were used: character error rate (CER) and Levenshtein distance. A lower score means a better performance. The answers with the lowest Levenshtein distance (i.e. the best score) were used for the calculation. See Table 8 for the scores.
Model CER Levenshtein
Claude 3.5 Sonnet 7.9% 0.018
Doubao Pro 20.9% 0.044
Ernie 4.0 34.4% 0.094
Gemini 1.5 Flash 34.7% 0.083
Gemini 1.5 Pro 15.3% 0.030
GPT4o 5.4% 0.009
GPT4o-mini 12.0% 0.023
SenseChat 54.4% 0.163
Aya 23 8B 96.6% 0.724
CLLM 6B 94.1% 0.859
CLLM 34B 23.4% 0.058
Yi 1.5 6B 99.0% 0.577
Yi 1.5 9B 97.2% 0.528
Yi 1.5 34B 79.6% 0.837
Gemma 2 2B 97.5% 0.524
Gemma 2 9B 73.0% 0.259
Gemma 2 27B 62.5% 0.201
Llama 3.1 8B 69.9% 0.270
Llama 3.1 70B 31.3% 0.086
Llama 3.1 405B 26.3% 0.074
Mistral Nemo 12B 59.8% 0.201
Qwen2 7B 97.3% 0.466
Qwen2 72B 74.0% 0.268
Rule Based 5.0% 0.009
@@ F.1 Translation Results Evaluation
A graphical user interface based on Label Studio was configured such that annotators (paid undergraduate students from Hong Kong) can highlight mistakes in the translated text for the following categories:

### v2

@@ 4.4 Linguistic and NLP Tasks
Most LLMs understand Cantonese fine. Most models performed well in Sentiment Analysis (GPT4o 79.7%, Llama 3.1 405B 78.8%), Translation (3-shot: GPT4o 98.3%, Qwen2 72B 96.6%), and Summarisation (Claude 3.5 Sonnet 92.7%, Gemma 2 9B 91.3%). Models that obtained lower scores are often due to task completion problems, e.g. failure to handle long input and problems with low-frequency/mixed-language tokens.
Proprietary and large open-weight models have good Cantonese lexical knowledge. The performance in translation and sentiment analysis is closely tied to the ability to determine the meaning of Cantonese-specific words that are not found or used differently in Mandarin. Most models also performed well in the Cantoense Character Selection sub-task (Canton. Char. in Table 2) under Orthographic Knowledge. It is noteworthy that despite good performance with proprietary models (73.1% - 88.5%) and some open-weight models (CLLM v0.5 34B and Qwen2 72B, both 76.9%), GPT4o struggled with Cantonese orthography (65.4%).
LLMs in general lack knowledge about Cantonese pronunciation. In the Grapheme-to-Phoneme (G2P) conversion task, all models performed far worse than the rule-based control (Visual Fonts v3.3, CER 0.8%), with the closest being GPT-4o (5.4%) and Claude 3.5 Sonnet (7.9%) as shown in Table 3. The appalling results from all tested language models reveal how linguistic knowledge is seriously under-represented. While it is expected that the G2P tasks will be significantly improved in newer/future models, actual linguistic tasks that involve sounds require more advanced knowledge about the language’s sound system. Most models struggled with tasks like judging homophone or rhyme pairs in Table 2, with GPT-4o being a notable exception (Homophone: 56.0%; Rhyming: 96.0%). Poor (close to chance level) performance in other models is not only due to the lack of G2P ability, a prerequisite for phonological reasoning, but also due to how Mandarin homophones partially influence this task. This will continue to be challenging for Cantonese due to limited specialised data.
LLMs in general do not have meta-linguistic knowledge represented in Cantonese. Although certain models, especially the Chinese proprietary models, performed well in the visual similarity task (SenseChat 70%, Doubao 70%, Ernie 75%) or orthographic reasoning (GPT4o 63.0%), the knowledge seems to have come from Simplified Chinese, thus their good performance is not transferred to Cantonese-specific items. This seems to be caused by insufficient descriptive knowledge about the structure and properties associated with the individual glyphs.
@@ 5 Conclusion
This paper presents HKCanto-Eval, the first comprehensive evaluation benchmark focusing on Hong Kong Cantonese, by comparing the Cantonese language support of 6 proprietary and 7 open-weight model families. Our findings indicate that while these models can understand Cantonese in various contexts, retrieve knowledge about Hong Kong, and address problems written in or about Cantonese to some extent, there are notable limitations. Most models, especially open-weight models in the 6–9B range, lack sufficient linguistic, cultural and professional knowledge in Cantonese and Hong Kong. Performance was particularly poor for questions requiring knowledge not commonly found in major online sources.
One area that we paid close attention to is the presence of metalinguistic knowledge in these models. There is concern that models showed Cantonese proficiency in linguistic and NLP tasks primarily through Mandarin. If their linguistic understanding is based solely on Mandarin, they may perform well on simpler tasks but struggle significantly with “false friends" between languages, as Mandarin knowledge becomes a hindrance. This benchmark introduces a novel perspective, focusing on Cantonese processing abilities beyond superficial slang and expressions. By requiring reasoning about sounds and characters specific to Cantonese, our benchmark provides a fairer judgement that credits models accurately capturing Cantonese phonology and orthography, while exposing those that appear competent in Cantonese but are heavily reliant on Mandarin.
@@ 6 Limitations & Future Directions
The current benchmark exhibits several limitations.
Inaccuracies in machine-translated materials: First, the use of machine translation introduces potential inaccuracies. While Gemini 1.5 Flash balances cost and quality, human-translated questions could provide a more reliable benchmark, albeit at a higher resource cost. The reliance on multiple-choice and text-based questions does not fully capture the capabilities required for practical LLM applications such as code generation and mathematical problem-solving, which demand coherent and contextual text generation. The dataset also lacks multi-modal data like image and audio, which is now supported by proprietary models and should be evaluated.
Biases in topic selection: The newly and manually created questions might contain biases and a lack of scalability and comprehensiveness. The cultural questions, predominantly created by colleagues and relatives of the authors, may introduce bias in cultural references and wordings, leading to an over-representation of certain perspectives while under-representing others, such as traditional practices. Political topics were also specifically excluded, due to political complications, limiting cultural representation. This can also be considered a reasonable compromise since many models (e.g. those from Chinese companies) are configured to censor these topics, and there is a risk that our accounts or IP addresses will be banned before we complete all the benchmarking tasks for this paper.
Lack of Crosslingual Evaluation: English translations for cross-lingual ability evaluation were also not included due to resource limitations. An additional comparison should be added to compare whether the same set of questions will be answered less satisfactorily when presented in English or Standard Written Chinese instead of Cantonese, in line with the evaluation done for Basque (Etxaniz et al., 2024a) and Mongolian and Tibetan (Zhang et al., 2025). We will leave this for future research.
Reliance on human evaluation: Human evaluation, while insightful, is not scalable. Automated and objective evaluation methods, such as LLM-as-a-judge or rule-based approaches, are necessary for efficient evaluation, but this is challenging due to the low-resource nature of Cantonese.
@@
Acknowledgments
Open-source models evaluated with computer resources offered under the categories of Trial Use and General Projects by the Research Institute for Information Technology, Kyushu University. Votee AI gratefully sponsored the usage fee. T.C.C is supported by the MEXT Initiative to Establish Next-Generation Novel Integrated Circuit Centers (X-NICS). The project is partially supported by funding from the Centre for Research on Linguistics and Language Studies (CRLLS), the Education University of Hong Kong.
Appendix
{CJK*}
UTF8bsmi
@@ Appendix A Questionable Practices in an Existing Work
Jiang et al. (2024) recently proposed a Cantonese evaluation dataset consisting of 5 datasets. The authors cited a HuggingFace organisation homepage in a footnote for their translation dataset but offered no further specifics. Yet, the dataset’s entries were not identifiable within the linked account. Nonetheless, the translation pairs between Cantonese and English are dubious. During the machine translation process, the less advanced models before GPT-3.5 tend to break English sentences into phrases by punctuation marks or connecting words such as "and" and "but". Then the phrases were translated individually and finally put together into one sentence in Cantonese. As a result, the translated texts were full of wrong wordings such as the following translation pair:
@@
The translation service broke down the original English sentences by punctuation and connectives (such as "and"). Then, it translated the broken-down phrases individually and joined them into a Cantonese/Chinese sentence, resulting in awkward and incorrect punctuation usage. The heavy use of poor machine translation of existing work offered little in terms of novel contributions or insights into the language.
@@ Appendix C Academic and Professional Dataset
    Insurance Intermediaries Qualifying Examination: The Insurance Authority in Hong Kong regulates and licenses all insurance intermediaries in Hong Kong and offers this exam for any person who wishes to carry out insurance activities.
@@
    Licensing Exam for Securities & Futures Intermediaries: An exam offered by the Hong Kong Securities and Investment Institute and recognised by the Securities and Futures Commission (SFC). It is a professional qualification exam for people who would like to work in the securities and investment industry in Hong Kong.
@@
The Academic Dataset subset was sourced from scanned copies of the HKDSE exam from the last 2 to 4 years. The details can be found in Table 5. The scanned PDFs were processed to extract text with the Google Gemini 1.5 Pro API. The extracted text was then corrected and filtered by the authors to remove questions that required information from other questions. Those requiring references to visual materials like pictures and maps were also removed but reserved for future evaluation of vision-enabled LLMs.
Although only multiple-choice questions were included in the dataset, more complicated questions that require complex reasoning and human judgment, such as Chinese listening and Liberal Studies exams were studied. Initial testing showed proprietary LLMs can easily achieve perfect scores in the Chinese listening exam when given the transcript, while the length of the transcript exceeded some open-weight LLMs’ context length. Further experiments were conducted with Google Gemini 1.5 Flash and Pro to take the audio file and written questions as input. Although the two models again achieved a perfect score, no other model supported audio input at the time of the experiment, so the task was excluded. For the now-defunct Liberal Studies, examinees were required to write a passage taking into account the provided reference materials and then develop their own point of view supported by examples from the candidates’ own knowledge. However, initial testing showed that none of the LLMs were able to include their own examples despite being explicitly requested in the prompt. These tasks were therefore not included in the final evaluation.
@@ Appendix E Linguistic Knowledge Dataset
This dataset comprises three sub-datasets with questions carefully designed with expert input from researchers specialising in different areas of Cantonese linguistics. The structure and questions of each dataset are outlined below.
@@ E.1 Phonological Knowledge Dataset
    MultiPron Resolution: A word that contains a character that can be pronounced in multiple ways (like the English word read can be pronounced as red or reed), and the task is to decide how this character should be pronounced. For example, if the given word is [Uncaptioned image][Uncaptioned image] and the question asks for a homophone for the 2nd character, given the list A [Uncaptioned image] B [Uncaptioned image] C [Uncaptioned image] D [Uncaptioned image], the answer is none of the above. The given character should be pronounced as ho4 and none of the answer (ho2, ho6, ho1, wo4) matches this pronunciation.
@@
    Poetry Rhyme: Given a poem, choose the correct combination of rhyming characters, e.g. in the poem [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]。 The last syllables are zeon6, lau4, muk6, lau4 in Cantonese, [Uncaptioned image] and [Uncaptioned image] are the two characters that rhyme in the poem when recited in Cantonese.
@@ E.3 Grapheme-to-Phoneme (G2P) Conversion Dataset
Score calculation: All acceptable variants have been listed as answers and all variants are considered equally good. This is to handle variant forms and ambiguous interpretations that may not be the standard, but native speakers of Cantonese accept. We will use the answer with the highest score for the subsequent calculation. Two metrics were used: character error rate (CER) and Levenshtein distance. A lower score means a better performance. The answers with the lowest Levenshtein distance (i.e. the best score) were used for the calculation.
@@ F.1 Translation Results Evaluation
A graphical user interface based on Label Studio was configured such that annotators (paid undergraduate students and teaching assistants from Hong Kong) can highlight mistakes in the translated text for the following categories: